OpenStudy (anonymous):

on a stretch of a road that climbs up a steep hill, there is a road sign that reads 9% grade. This means that the slope of the road is 9/100. over a certain timeyour elavation increases by 200 feet. What is the horizontal change in your position?

OpenStudy (anonymous):

times top and bottom by 200 200/200*9/100 =1800/20000 divide 9 =200/2222 answer =2222 feet